- 02-08-18Netwerk onderhoud
- 01-08-18Netwerk onderhoud
- 31-07-18Netwerk onderhoud
- 20-07-18Periodiek Onderhoud Stroomvoorziening BIT-2BCD
- 19-07-18Periodiek Onderhoud Stroomvoorziening BIT-2BCD
- 28-06-18Software-upgrade core routers
- 27-06-18Onderhoud UPS'en BIT-1
- 26-06-18Periodiek onderhoud UPS'en BIT-2D
- 22-06-18Onderhoud koeling BIT-1
- 22-06-18Periodiek onderhoud UPS'en BIT-2B Feed-B
Outage fiber BIT-1 <-> TC2
Around 14.39h we've noticed a disruption in the services provided on the fiber between BIT-1 and Telecity 2 in Amsterdam. Chances are that you may have experienced connection issues. Customers that have a separate 'wave' between BIT-1 and Telecity could still experience downtime for their connection.
Engineers are working hard to find the root cause and possible solution.
Updates will be posted on this page.
If you have any questions or comments please don't hesitate to contact us by telephone at +31(0)318 648 688 or by email at firstname.lastname@example.org.
Update placed on 31-10-2017, 15.07h:
The fiber connection between BIT-1 en Telecity 2 is down. Traffic is diverted towards other links. Custromers with a wave between these location are experiencing downtime.
Update placed on 31-10-2017, 16.20h:
Our supllier confirms that the fibercut is somewhere between Doorn and Amsterdam and is investigating the situation. No estimate on recovery time can be given at this moment.
Update placed on 31-10-2017, 20.20h:
Our supplier has determined the approximate location of the fiber cut, which seems to be near Utrecht. Engineers are enroute to the suspected break location to investigate damage.
Update placed on 31-10-2017, 22.45h:
The supplier has located the damage. It appears a construction crew has bored through the fiber. The engineers are still assessing the damage at this time.
Update placed on 01-11-2017, 00.15h:
The onsite engineers of the supplier have assessed the damage. Multiple repairs are required to resolve the fibercut, which will take some time. Replacement material that is required for these repairs is underway. No estimate on recovery time can be given at this moment.
Update placed on 01-11-2017, 01.10h:
Materials to create a temporary solution have arrived on-site. Excavators are still working on the temporary solution. Because of this the new ducts cannot be placed yet. No estimate on recovery time can be given at this moment.
Update placed on 01-11-2017, 06.30h:
A number of cranes have dug up the damaged fiberduct. A large number of fiber pairs need to be restored. A number of fiber splicing crews is at the site working on these fibers. Our supplier expects the fibers to be restored at 10h.
BIT's fiber seems to be operational again, but due to the increased risk of disruptions during the restoration works of other fibers in this duct traffic is still rerouted via other connections until we have a confirmation from our supplier that all restoration works have been completed.
Update placed on 01-11-2017, 07.15h:
Restoration works are progressing faster than expected. The new time to fix has been set at 9h. If works continue to go faster than expected all services can be restored earlier than that.
Update placed on 01-11-2017, 08.45h:
Restoration works have been completed. Our fiber connection has been up and stable for over two hours and is now in production again.
The incident has been resolved. If you have any questions regarding this incident, please contact the Customer Care department by phone on +31 (0)318 648 688 or by email on email@example.com.